How much Ca3(PO4)2(s) could be produced in an industrial process if 55.00 g of CaCl2 in solution reacted completely with sufficient Na3(PO4)(aq)?


1
Expert's answer
2021-03-26T05:39:13-0400

According to the reaction:

3CaCl2+ 2Na3(PO4) = Ca3(PO4)2 + 6NaCl

From here, the mass of Ca3(PO4)2 equals:

m(Ca3(PO4)2) = m(CaCl2) × Mr(Ca3(PO4)2) / (Mr(CaCl2) × 3) = 55.00 g × 310 g/mol / (111 g/mol × 3) = 51.2 g


Answer: 51.2 g